The museum charges different admission prices for adults and kids. Admission to the museum for 28 kids and 5 adults cost $284. Admission for 40 kids and 10 adults costs $440. What is the admission for one kid to the museum?
Answer by rumessa(13)

one kid =x
one adult = y
28x +5y = 284
40x + 10y = 440
2(28x + 5y = 284) 56x + 10y = 568
1(40x + 10y = 440) 40x + 10y = 440
- - -
____________________
16x = 128
x = 8
one kid admission = $8